OpenStudy (anonymous):

suppose that integral from 1 to x f(t) dt=5x^2+6x-2. Find f(x)

OpenStudy (amistre64):

are the little fs derivatives of big Fs ?

OpenStudy (amistre64):

\[\int_{A_x}^{B_x}f(t)dt=F(B_x)-F(A_x)\] \[D_x[F(B_x)-F(A_x)]=f(B_x)B'-f(A_x)A'\] \[D_x[F(B_x)-F(A_x)]=f(x)-f(1)0=f(x)\]

OpenStudy (amistre64):

looks to me that f(x) = 5x^2+6x-2

OpenStudy (anonymous):

thanks